10. The probability that automobile sales will increase next month (A) is estimated to be 0.40. The probability that the sale of replacement parts will increase (R) is estimated to be 0.50. The probability that both industries will experience an increase in sales is estimated to be 0.10. What is the probability that (a) automobile sales have increased during the month given that there is information that replacement parts sales have increased, (b) replacement parts sales have increased given information that automobile sales have increased during the month? Are the two events dependent or independent?
Answer #1

P(A) = 0.4

P(R) = 0.5

P(A and R) = 0.1

a) P(A | R) = P(A and R) / P(R) = 0.1 / 0.5 = 0.2

b) P(R | A) = P(A and R) / P(A) = 0.1 / 0.4 = 0.25

The two events are not independent. Because P(A | R) is not equal to P(A)